# Hypersonic flow onto a large curved wedge and the dissipation of shock wave

### v1

@@ 4.1.1. $(\xi,\eta)$ Coordinates
Clearly, $P_{w}P$ is perpendicular to $\mathsf{W}$ at $P_{w}$. Here and in sequel, for any point $P$, the subscript $w$ stands for its projection on $\mathsf{W}$, and the superscript $s$ stands for the intersection of $\mathsf{S}$ and $P_{w}P$. Denote the intersection of $P_{w}P$ and $\mathsf{S}$ by $P^{s}$. It is obvious to have the following property about the transformation.
@@ 4.1.2. Thickness of the shock layer
    with $M_{0,1}$ being constants depending only on $\underline{f}$ and $\bar{f}$.
@@
where $\vec{U}_{s}\in\mathcal{C}_{s}$ and $\xi\in\mathbb{R}^{+}$. By a perturbed discussion, we could prove (4.1.3) by (4.1.6).
@@
Next, by conservation of mass (see Figure 5),we have
@@
For the left-hand side of the above equality, since by Lemma 3.5 the slope $\phi^{\prime}(x)=\phi^{\prime}(\vec{U},\epsilon)$ of the shock-front is control by $\underline{\phi}$ and $\bar{\phi}$, we have on the one hand
@@ 4.1.3. A priori estimates related to the characteristics in the shock layer
In order to estimate the length of the characteristics between the shock $\mathsf{S}$ and the wall $\mathsf{W}$, we consider a standardised domain $\Omega$. It has a wall boundary $\mathsf{W}$ and a shock boundary $\mathsf{S}$. Let $P_{w}\in\mathsf{W}$ and $P^{s}\in\mathsf{S}$. We employ a arclength parameter $\xi$ on $\mathsf{W}$ with $\xi(P_{w})=0$ and $\left|P_{w}P^{s}\right|=1$. $\eta$ is the distance function with respect to $\mathsf{W}$. Assume a $C^{-}$ characteristic starting from $P^{s}$ intersects with a $C^{+}$ characteristic starting from $P_{w}$ at $Q$. We remind that $Q_{w}$ is the projection of $Q$ on the wall and $Q^{s}$ is the intersection of $QQ_{w}$ and $\mathsf{S}$. Then we have the following characteristic estimates result.
@@
(1) the length of $\widehat{R\bar{R}_{1}},\widehat{\bar{R}_{1}\underline{R}_{2}},\widehat{R\underline{R}_{1}}$ and $\widehat{\underline{R}_{1}\bar{R}_{2}}$ are less than some constant $M_{l}$, which is only depending on $\omega_{0}$;
@@
In view of above, as illustrated in Figure 7, for the constant states $\vec{U}_{s}$, the standardised domain becomes a domain bounded by a pair of parallel lines $\mathsf{W}$ and $\mathsf{S}$. The characteristics become two groups of parallel lines have an angle $\omega(\vec{U}_{s})$ with $\mathsf{W}$ and $\mathsf{S}$. Thus in view of Lemma 3.3 and Lemma 3.4, the wave pattern in Figure 6 is only a perturbation of that in Figure 7. The conclusion is obvious.
@@
which is corresponding to case where $\vec{U}(P)\equiv\vec{U}_{s}(P_{w})$ and $\epsilon=0$. The other cases for the first part can be discussed similar.
@@ 4.2.1. Inductive estimates and the local solution
There exists a positive constant $\epsilon_{O}$, such that, for any $\epsilon\leq\epsilon_{O}$, Problem 2 admits a local solution with the inductive estimates: (4.2.2) and (4.2.3) in $\xi\leq\xi_{\epsilon}$, where $\xi_{\epsilon}$ is a constant depending on $\epsilon$.
@@
By in [53, Theorem 3.1.1.], we can direct establish local existence of $C^{1}$ solution for Problem 2. The solution $\vec{U}$ and $\vec{U}_{s}$ obey the algebra equations
@@
Furthermore, if it holds (4.2.2) in $\Omega_{0}$ for $\xi(x,y)\in[0,\xi_{0}]$, we have
@@
For $(u,v)\in\mathcal{C}_{s}^{\delta_{\omega}}$, it follows from Lemma 3.5 that the vectors $(\cos\alpha,\sin\alpha)$ and $(\cos\beta,\sin\beta)$ have an angle, that is $2\omega$, between $[2\omega_{0},\pi-2\omega_{0}]$. So, for any smooth function $\psi$, we have the following derivative estimates
@@
In view of above two, the state $(u,v)(P)$ must locates around the insection of the shock polar (2.1.2) and the line (4.2.10).
In addition, let $\vec{U}_{\epsilon}^{*}=(u_{\epsilon}^{*},v_{\epsilon}^{*})$ be the unique supersonic states given by
@@ 4.2.2. Finite Growth Estimate
By Theorem 4.4, we know that it holds (4.2.2) and (4.2.3) for $\xi\leq\xi_{\epsilon}$. We will show that if the solution $(u,v)$ satisfies (4.2.2) and (4.2.3) for $\xi\leq\xi_{0}$ with $\xi_{0}>0$, then there exists $\Delta$ depending on $\xi_{0}$ such that the solution can be extended to $\xi\leq\xi_{0}+\Delta$ with (4.2.2) and (4.2.3). As illustrating in Figure 8, with given smooth data on $P_{w}P^{s}=\left\{P|\xi(P)=\xi_{0},~{}0\leq\eta(P)\leq\eta(P^{s})\right\}$ and the slipping boundary condition on the wedge $\mathsf{W}$, we consider the standardised free boundary value problem in $\xi\geq\xi_{0}$. We have the following finite growth estimates.
@@
To see the impact of $\epsilon$, we employ a standard scaling transformation discussion. Apply the coordinates transformation
@@
Note that it holds (4.2.19) on $P_{w}P^{s}$. By continuity of the solution as well as its first derivative, we know that (4.2.20) hold in a neighbourhood of $P_{w}P^{s}$. We first prove that (4.2.20) could not fail in $\Omega_{1}$ under the assumption (4.2.19) on $P_{w}P^{s}$. Assume to the contrary that there exists a point $P^{\prime}\in\Omega_{1}$ such that (|¯$\left(\left|\bar{\partial}^{+}c\right|+\left|\bar{\partial}^{-}c\right|\right)(P^{\prime})M_{\omega}=1$, (|¯$\left(\left|\bar{\partial}^{+}c\right|+\left|\bar{\partial}^{-}c\right|\right)(Q)M_{\omega}<1$ and $\vec{U}(Q)\in\mathcal{C}_{s}^{\delta_{0}}$ for $Q\in\Omega_{1}$ with $\xi(Q)<\xi(P^{\prime})$.We go on to deduce a contradiction.
@@
If it holds (4.2.2) and (4.2.3) for $\xi\leq\xi_{0}$, then there exists $\epsilon_{B}\leq\epsilon_{F}$, such that in the domain
@@
By the uniformly continuity of $\vec{U}_{s}(\xi)$ and the decay (1.0.7), for sufficiently small $\epsilon_{L}$, we can show (4.2.28) for $\xi\in[(1-\epsilon M_{0}\frac{8}{\tan\omega_{0}})\xi_{0},\xi_{0}]$ and $\xi_{0}\in\mathbb{R}^{+}$. Thus (4.2.28) is proved.
@@
Next, we intend to show the inductive estimates (4.2.2) and (4.2.3) for $\xi\leq\xi_{0}+\frac{1}{8}\tan\omega_{0}\left|P^{s}P_{w}\right|(\xi_{0})$. First, we prove (4.2.2) in the following
@@
By the above theorem, we get (4.2.2) in this region. Then, by Theorem 4.7, the assumptions of Theorem 4.6 is valid, we have (4.2.8). ∎
@@ 4.3. Asymptotic behaviour
for $\xi(P)\geq X$. As indicated in Figure 10, for any point $P\in\Omega$, denote by $C_{1}^{-}$ the backward $C^{-}$ characteristic passing through it. This characteristic intersects the shock front S at $\bar{P}_{1}$. Then, we denote by $C_{1}^{+}$ the backward $C^{+}$ characteristic starting from $\bar{P}_{1}$ and $\underline{P}_{1}$ the intersection of $C_{1}^{+}$ and W. Such a procedure will continue and finally approximate to the origin.
@@ 4.4. Higher order properties
Thus, for fixed $\epsilon>0$, the characteristic finally approximate the origin $\xi=0$, otherwise it leads to a contradiction. We obtain a series $\bar{P}_{n}$ on the shock S and a series $\underline{P}_{n}$ on the wedge boundary W with the $C^{\pm}_{n}$ characteristics.
@@ 5.1. Some sign results along the shock polar
We can see in the above proof that the lower positive bound of $v$ plays an important role to ensure the valid of (5.1.1). If the wedge boundary approaches to flat, this assumption on $v$ will be invalid. It makes the estimates similar to (5.1.1) be come much more complex. For this case, we have the following
@@
$\frac{U}{Y}\rightarrow\mbox{the roots of}\quad{x^{2}-12x+3=0},$
@@
$\frac{U}{Y}\rightarrow\mbox{the roots of}\quad{x^{3}-7x^{2}+11x-5=0}.$
@@
for $(u,v)$ around $(\bar{q},0)$. Then, as a result, (5.1.3) would be proved by combining (5.1.14) and (5.1.15). To do it, first of all, we have $\sin(\frac{\pi}{2}+k-\alpha)=0=\sin(n-\alpha)$. Then for $(u,v)$ around $(\bar{q},0)$, $c=\sqrt{\frac{\bar{q}^{2}}{2}-\frac{u^{2}+v^{2}}{2}}$ is small. So does the quantity $\omega=\arcsin\frac{c}{\sqrt{u^{2}+v^{2}}}$. Due to $\beta=\alpha-2\omega$, we have by equivalent infinitesimals the following
@@
So $F(u,v)=\rm{constant}$ along the $C^{-}$ characteristic. By the equation ¯$\bar{\partial}^{-}u+\lambda_{+}\bar{\partial}^{-}v=0$, we have
@@
By the above equation, contour lines $F(u,v)=\rm{constant}$ are the integral curves defined by the vector fields $(\lambda_{+},-1)$ or $(\sin\alpha,-\cos\alpha)$. The conclusions follow by a comparison of $(\sin\alpha,-\cos\alpha)$ with the following vectors:
  • the tangent vector of circle $q=\mbox{constant.}$.: $(\sin\tau,-\cos\tau)$;
  • the inner normal vector of circle $q=\mbox{constant.}$.: $(-\cos\tau,-\sin\tau)$;
@@
In a similar way, the second inequality of (5.1.22) indicates that $q$ increases in $(\sin\alpha,-\cos\alpha)$ direction. The third inequality of (5.1.22) indicates that $(\sin\alpha,-\cos\alpha)$ inserts the domain bounded by the shock polar. Since the curve $F(u,v)=\rm{constant}$ is generated by the vector fields $(\sin\alpha,-\cos\alpha)$, combine above first two conclusions, along these curves, if $\tau$ decreases, then $q$ increases and so $c$ decreases.
Meanwhile, in view of the third conclusion of (5.1.22), the curve must insert the shock polar provided $c$ decreases. Remind that, along each $C^{-}$ characteristic, the states must locate on the curve $F(u,v)=\rm{constant}$, we can summarise above and arrive at the first conclusion of this theorem.
@@ 5.2. Second initial value condition
Under the assumption of Case B, for some positive $x_{0}>0$ with $\xi_{0}=\xi(x_{0},f(x_{0}))$, we have
@@ 5.3. Global solution
As illustrating in Figure 12, we consider the pattern on $\mathsf{S}$ locally. Let $\bar{P}_{1}$ and $\bar{P}_{2}$ on $\mathsf{S}$ be two adjacent points connecting the $C^{\pm}$ characteristics with the intersection $P$. By ¯±$\bar{\partial}^{\pm}c<0$, we have $c|_{\bar{P}_{1}}>c|_{P}>c|_{\bar{P}_{2}}$ locally. This indicates $q|_{\bar{P}_{1}}>q|_{\bar{P}_{2}}$. So the strength of the shock is weaken locally along the shock $\mathsf{S}$ as $\xi$ increasing. Moreover, $s$ is decreasing. Thus, the shock is convex locally. We complete the proof. ∎
Then, to show the global estimates, we prove a continuity lemma. As illustrated in Figure 8, we assume that the $C^{-}$ characteristic connects the point $P^{s}$ on $\mathsf{S}$ with the point $\underline{P}_{1}$ on $\mathsf{W}$ and the $C^{+}$ characteristic connects the point $P_{w}$ on $\mathsf{W}$ with the point $\bar{P}_{1}$ on $\mathsf{S}$. Therefore, the domain is divided into $\Omega_{1}$, $\bar{\Omega}_{1}$ and $\underline{\Omega}_{1}$. We have
@@
In view of the characteristic decomposition (2.2.9), we see that ¯±$\bar{\partial}^{\pm}c$ can not change sign along the characteristic. So ¯±$\bar{\partial}^{\pm}c|_{P^{s}P_{w}}<0$ indicates ¯$\bar{\partial}^{-}c<0$ in $\Omega_{1}\cup\bar{\Omega}_{1}$ and ¯$\bar{\partial}^{+}c<0$ in $\Omega_{1}\cup\underline{\Omega}_{1}$. On the shock $\mathsf{S}$, ¯$\bar{\partial}^{-}c<0$, (2.4.8) and (5.2.5) yield ¯$\bar{\partial}^{+}c<0$ on $\mathsf{S}$ provided that (5.3.4) holds. Since we have (5.3.1), (5.3.4) holds around $P^{s}$ on $\mathsf{S}$. By Lemma 5.8, ¯±$\bar{\partial}^{\pm}c<0$ guarantee the convexity of the shock $\mathsf{S}$. On the one hand, the convexity makes $u$ increase along $\mathsf{S}$. So $u|_{\mathsf{S}\cap\left\{\xi\geq\xi(P^{s})\right\}}\geq u(P^{s})$. On the other hand, (2.2.8) indicates $c\bar{\partial}^{+}\tau=-\frac{\sin(2\omega)}{2\kappa}\bar{\partial}^{+}c>0$ provided ¯$\bar{\partial}^{+}c<0$. So $\tau|_{\mathsf{S}\cap\left\{\xi\geq\xi(P^{s})\right\}}\geq\tau|_{P^{s}P_{w}}>0$. These two reinforce the condition (5.3.4) along $\mathsf{S}$. Therefore, (5.3.4) must hold along the shock $\mathsf{S}$ between $P^{s}$ and $\bar{P}_{1}$. We have ¯$\bar{\partial}^{+}c<0$ in $\bar{\Omega}_{1}$. On the other hand, ¯$\bar{\partial}^{+}c<0$, convex condition $\kappa_{w}\leq 0$ and (2.3.2) yield ¯$\bar{\partial}^{-}c<0$ on $\mathsf{W}$. Finally, we get ¯$\bar{\partial}^{-}c<0$ in $\underline{\Omega}_{1}$. Since (5.3.3) and $\tau|_{\mathsf{W}}\geq 0$, (5.3.5) and $c\bar{\partial}^{-}\tau=\frac{\sin(2\omega)}{2\kappa}\bar{\partial}^{-}c<0$ in (2.2.8) lead to (5.3.6). ∎

### v2

@@ 4.1.1. $(\xi,\eta)$ Coordinates
Clearly, $P_{w}P$ is perpendicular to $\mathsf{W}$ at $P_{w}$. Here and in the sequel, for any point $P$, the subscript $w$ stands for the projection of $P$ on $\mathsf{W}$, and the superscript $s$ stands for the intersection of $\mathsf{S}$ and $P_{w}P$. Denote the intersection of $P_{w}P$ and $\mathsf{S}$ by $P^{s}$. It is obvious that the transformation has the following property.
@@ 4.1.2. Thickness of the shock layer
    where $M_{0,1}$ are constants depending only on $\underline{f}$ and $\bar{f}$.
@@
where $\vec{U}_{s}\in\mathcal{C}_{s}$ and $\xi\in\mathbb{R}^{+}$. By a perturbation discussion, we could prove (4.1.3) by (4.1.6).
@@
Next, by the conservation of mass (see Figure 5),we have
@@
For the left-hand side of the above equality, since by Lemma 3.5 the slope $\phi^{\prime}(x)=\phi^{\prime}(\vec{U},\epsilon)$ of the shock-front is controlled by $\underline{\phi}$ and $\bar{\phi}$, we have on the one hand
@@ 4.1.3. A priori estimates related to the characteristics in the shock layer
In order to estimate the length of the characteristics between the shock $\mathsf{S}$ and the wall $\mathsf{W}$, we consider a standardised domain $\Omega$. It has a wall boundary $\mathsf{W}$ and a shock boundary $\mathsf{S}$. Let $P_{w}\in\mathsf{W}$ and $P^{s}\in\mathsf{S}$. We employ an arclength parameter $\xi$ on $\mathsf{W}$ with $\xi(P_{w})=0$ and $\left|P_{w}P^{s}\right|=1$. $\eta$ is the distance function with respect to $\mathsf{W}$. Assume a $C^{-}$ characteristic starting from $P^{s}$ intersects with a $C^{+}$ characteristic starting from $P_{w}$ at $Q$. We remind that $Q_{w}$ is the projection of $Q$ on the wall and $Q^{s}$ is the intersection of $QQ_{w}$ and $\mathsf{S}$. Then we have the following characteristic estimates result.
@@
(1) the length of $\widehat{R\bar{R}_{1}},\widehat{\bar{R}_{1}\underline{R}_{2}},\widehat{R\underline{R}_{1}}$ and $\widehat{\underline{R}_{1}\bar{R}_{2}}$ are each less than some constant $M_{l}$, which is only depending on $\omega_{0}$;
@@
In view of above, as illustrated in Figure 7, for the constant states $\vec{U}_{s}$, the standardised domain becomes a domain bounded by a pair of parallel lines $\mathsf{W}$ and $\mathsf{S}$. The characteristics become two groups of parallel lines having an angle $\omega(\vec{U}_{s})$ with $\mathsf{W}$ and $\mathsf{S}$. Thus in view of Lemma 3.3 and Lemma 3.4, the wave pattern in Figure 6 is only a perturbation of that in Figure 7. The conclusion is obvious.
@@
which is corresponding to case where $\vec{U}(P)\equiv\vec{U}_{s}(P_{w})$ and $\epsilon=0$. The other cases for the first part can be discussed similarly.
@@ 4.2.1. Inductive estimates and the local solution
There exists a positive constant $\epsilon_{O}$, such that, for any $\epsilon\leq\epsilon_{O}$, Problem 2 admits a local solution with the inductive estimates (4.2.2) and (4.2.3) in $\xi\leq\xi_{\epsilon}$, where $\xi_{\epsilon}$ is a constant depending on $\epsilon$.
@@
By [53, Theorem 3.1.1.], we can directly establish local existence of $C^{1}$ solution for Problem 2. The solution $\vec{U}$ and $\vec{U}_{s}$ obey the algebra equations
@@
Furthermore, if (4.2.2) holds in $\Omega_{0}$ for $\xi(x,y)\in[0,\xi_{0}]$, we have
@@
For $(u,v)\in\mathcal{C}_{s}^{\delta_{\omega}}$, it follows from Lemma 3.5 that the vectors $(\cos\alpha,\sin\alpha)$ and $(\cos\beta,\sin\beta)$ have an angle, which is $2\omega$, in the interval $[2\omega_{0},\pi-2\omega_{0}]$. So, for any smooth function $\psi$, we have the following derivative estimates
@@
In view of above two, the state $(u,v)(P)$ must locates around the intersection of the shock polar (2.1.2) and the line (4.2.10).
In addition, let $\vec{U}_{\epsilon}^{*}=(u_{\epsilon}^{*},v_{\epsilon}^{*})$ be the unique supersonic state given by
@@ 4.2.2. Finite Growth Estimate
By Theorem 4.4, we know that it holds (4.2.2) and (4.2.3) for $\xi\leq\xi_{\epsilon}$. We will show that if the solution $(u,v)$ satisfies (4.2.2) and (4.2.3) for $\xi\leq\xi_{0}$ with $\xi_{0}>0$, then there exists $\Delta$ depending on $\xi_{0}$ such that the solution can be extended to $\xi\leq\xi_{0}+\Delta$ with (4.2.2) and (4.2.3). As illustrated in Figure 8, with given smooth data on $P_{w}P^{s}=\left\{P|\xi(P)=\xi_{0},~{}0\leq\eta(P)\leq\eta(P^{s})\right\}$ and the slipping boundary condition on the wedge $\mathsf{W}$, we consider the standardised free boundary value problem in $\xi\geq\xi_{0}$. We have the following finite growth estimates.
@@
To see the impact of $\epsilon$, we employ a standard scaling transformation. Apply the coordinates transformation
@@
Note that it holds (4.2.19) on $P_{w}P^{s}$. By continuity of the solution as well as its first derivative, we know that (4.2.20) holds in a neighbourhood of $P_{w}P^{s}$. We first prove that (4.2.20) could not fail in $\Omega_{1}$ under the assumption (4.2.19) on $P_{w}P^{s}$. Assume to the contrary that there exists a point $P^{\prime}\in\Omega_{1}$ such that (|¯$\left(\left|\bar{\partial}^{+}c\right|+\left|\bar{\partial}^{-}c\right|\right)(P^{\prime})M_{\omega}=1$, (|¯$\left(\left|\bar{\partial}^{+}c\right|+\left|\bar{\partial}^{-}c\right|\right)(Q)M_{\omega}<1$ and $\vec{U}(Q)\in\mathcal{C}_{s}^{\delta_{0}}$ for $Q\in\Omega_{1}$ with $\xi(Q)<\xi(P^{\prime})$.We go on to deduce a contradiction.
@@
If (4.2.2) and (4.2.3) hold for $\xi\leq\xi_{0}$, then there exists $\epsilon_{B}\leq\epsilon_{F}$, such that in the domain
@@
By the uniform continuity of $\vec{U}_{s}(\xi)$ and the decay (1.0.7), for sufficiently small $\epsilon_{L}$, we can show (4.2.28) for $\xi\in[(1-\epsilon M_{0}\frac{8}{\tan\omega_{0}})\xi_{0},\xi_{0}]$ and $\xi_{0}\in\mathbb{R}^{+}$. Thus, (4.2.28) is proved.
@@
Next, we intend to show the inductive estimates (4.2.2) and (4.2.3) for
$\xi\leq\xi_{0}+\frac{1}{8}\tan\omega_{0}\left|P^{s}P_{w}\right|(\xi_{0}).$
First, we prove (4.2.2) in the following
@@
By the above theorem, we get (4.2.2) in this region. Then, by Theorem 4.7, the assumptions of Theorem 4.6 are valid, we have (4.2.8). ∎
@@ 4.3. Asymptotic behaviour
for $\xi(P)\geq X$. As indicated in Figure 10, for any point $P\in\Omega$, denote by $C_{1}^{-}$ the backward $C^{-}$ characteristic passing through it. This characteristic intersects the shock front S at $\bar{P}_{1}$. Then, we denote by $C_{1}^{+}$ the backward $C^{+}$ characteristic starting from $\bar{P}_{1}$ and $\underline{P}_{1}$ the intersection of $C_{1}^{+}$ and W. This process continues and eventually approaches the origin.
@@ 4.4. Higher order properties
Thus, for fixed $\epsilon>0$, the characteristic finally approaches the origin $\xi=0$, otherwise it leads to a contradiction. We obtain a series $\bar{P}_{n}$ on the shock S and a series $\underline{P}_{n}$ on the wedge boundary W with the $C^{\pm}_{n}$ characteristics.
@@ 5.1. Some sign results along the shock polar
We can see in the above proof that the lower positive bound of $v$ plays an important role to ensure the validity of (5.1.1). If the wedge boundary approaches a flat, this assumption on $v$ will be invalid. It makes the estimates similar to (5.1.1) become much more complex. For this case, we have the following
@@
$\frac{U}{Y}\rightarrow\mbox{a root of}\quad{x^{2}-12x+3=0},$
@@
$\frac{U}{Y}\rightarrow\mbox{a root of}\quad{x^{3}-7x^{2}+11x-5=0}.$
@@
for $(u,v)$ around $(\bar{q},0)$. Then, as a result, (5.1.3) would be proved by combining (5.1.14) and (5.1.15). To do it, first of all, we have $\lim\limits_{(u,v)\rightarrow(\bar{q},0)}\sin(n-\alpha)=0$. Then for $(u,v)$ around $(\bar{q},0)$, $c=\sqrt{\frac{\bar{q}^{2}}{2}-\frac{u^{2}+v^{2}}{2}}$ is small. So does the quantity $\omega=\arcsin\frac{c}{\sqrt{u^{2}+v^{2}}}$. Due to $\beta=\alpha-2\omega$, we have by equivalent infinitesimals the following
@@
So
$F(u,v)=\rm{constant}$
along the $C^{-}$ characteristic. By the equation ¯$\bar{\partial}^{-}u+\lambda_{+}\bar{\partial}^{-}v=0$, we have
@@
By the above equation, contour lines
$F(u,v)=\rm{constant}$
are the integral curves defined by the vector fields $(\lambda_{+},-1)$ or $(\sin\alpha,-\cos\alpha)$. The conclusions follow by a comparison of $(\sin\alpha,-\cos\alpha)$ with the following vectors:
  • the tangent vector of circle $q=\mbox{constant}$: $(\sin\tau,-\cos\tau)$;
  • the inner normal vector of circle $q=\mbox{constant}$: $(-\cos\tau,-\sin\tau)$;
@@
In a similar way, the second inequality of (5.1.22) indicates that $q$ increases in $(\sin\alpha,-\cos\alpha)$ direction. The third inequality of (5.1.22) indicates that $(\sin\alpha,-\cos\alpha)$ inserts the domain bounded by the shock polar. Since the curve $F(u,v)=\rm{constant}$ is generated by the vector fields $(\sin\alpha,-\cos\alpha)$, combine the first two conclusions above, along these curves, if $\tau$ decreases, then $q$ increases and, consequently, $c$ decreases.
Meanwhile, in view of the third conclusion of (5.1.22), the curve must insert the shock polar provided $c$ decreases. Recall that along each $C^{-}$ characteristic, the states must lie on the curve $F(u,v)=\rm{constant}$. Summarising the above, we arrive at the first conclusion of this theorem.
@@ 5.2. Second initial value condition
Under the assumption of Case B, for some $x_{0}>0$ with $\xi_{0}=\xi(x_{0},f(x_{0}))$, we have
@@ 5.3. Global solution
As illustrated in Figure 12, we consider the pattern on $\mathsf{S}$ locally. Let $\bar{P}_{1}$ and $\bar{P}_{2}$ on $\mathsf{S}$ be two adjacent points connecting the $C^{\pm}$ characteristics with the intersection $P$. By ¯±$\bar{\partial}^{\pm}c<0$, we have $c|_{\bar{P}_{1}}>c|_{P}>c|_{\bar{P}_{2}}$ locally. This indicates $q|_{\bar{P}_{1}}>q|_{\bar{P}_{2}}$. So the strength of the shock is weakened locally along the shock $\mathsf{S}$ as $\xi$ increases. Moreover, $s$ is decreasing. Thus, the shock is convex locally. We complete the proof. ∎
Then, to show the global estimates, we prove the following continuity lemma. As illustrated in Figure 8, we assume that the $C^{-}$ characteristic connects $P^{s}$ on $\mathsf{S}$ with $\underline{P}_{1}$ on $\mathsf{W}$ and the $C^{+}$ characteristic connects $P_{w}$ on $\mathsf{W}$ with $\bar{P}_{1}$ on $\mathsf{S}$. Therefore, the domain is divided into $\Omega_{1}$, $\bar{\Omega}_{1}$ and $\underline{\Omega}_{1}$. We have
@@
In view of the characteristic decomposition (2.2.9), we see that ¯±$\bar{\partial}^{\pm}c$ can not change sign along the characteristic. So ¯±$\bar{\partial}^{\pm}c|_{P^{s}P_{w}}<0$ indicates ¯$\bar{\partial}^{-}c<0$ in $\Omega_{1}\cup\bar{\Omega}_{1}$ and ¯$\bar{\partial}^{+}c<0$ in $\Omega_{1}\cup\underline{\Omega}_{1}$. On the shock $\mathsf{S}$, ¯$\bar{\partial}^{-}c<0$, (2.4.8) and (5.2.5) imply ¯$\bar{\partial}^{+}c<0$ on $\mathsf{S}$ provided that (5.3.4) holds. Since we have (5.3.1), (5.3.4) holds around $P^{s}$ on $\mathsf{S}$. By Lemma 5.8, ¯±$\bar{\partial}^{\pm}c<0$ guarantee the convexity of the shock $\mathsf{S}$. On the one hand, the convexity makes $u$ increase along $\mathsf{S}$. So $u|_{\mathsf{S}\cap\left\{\xi\geq\xi(P^{s})\right\}}\geq u(P^{s})$. On the other hand, (2.2.8) indicates $c\bar{\partial}^{+}\tau=-\frac{\sin(2\omega)}{2\kappa}\bar{\partial}^{+}c>0$ provided ¯$\bar{\partial}^{+}c<0$. So $\tau|_{\mathsf{S}\cap\left\{\xi\geq\xi(P^{s})\right\}}\geq\tau|_{P^{s}P_{w}}>0$. These two reinforce the condition (5.3.4) along $\mathsf{S}$. Therefore, (5.3.4) must hold along the shock $\mathsf{S}$ between $P^{s}$ and $\bar{P}_{1}$. We have ¯$\bar{\partial}^{+}c<0$ in $\bar{\Omega}_{1}$. On the other hand, ¯$\bar{\partial}^{+}c<0$, convex condition $\kappa_{w}\leq 0$ and (2.3.2) imply ¯$\bar{\partial}^{-}c<0$ on $\mathsf{W}$. Finally, we get ¯$\bar{\partial}^{-}c<0$ in $\underline{\Omega}_{1}$. Since (5.3.3) and $\tau|_{\mathsf{W}}\geq 0$, (5.3.5) and $c\bar{\partial}^{-}\tau=\frac{\sin(2\omega)}{2\kappa}\bar{\partial}^{-}c<0$ in (2.2.8) lead to (5.3.6). ∎